# PAFT: Prompt-Agnostic Fine-Tuning

### v1

@@ Abstract
While Large Language Models (LLMs) adapt well to downstream tasks after fine-tuning, this adaptability often compromises prompt robustness, as even minor prompt variations can significantly degrade performance. To address this, we propose prompt-agnostic fine-tuning (PAFT), a simple yet effective approach that dynamically adjusts prompts during fine-tuning. This encourages the model to learn underlying task principles rather than overfitting to specific prompt formulations. PAFT operates in two stages: First, a diverse set of meaningful, synthetic candidate prompts is constructed. Second, during fine-tuning, prompts are randomly sampled from this set to create dynamic training inputs. Extensive experiments across diverse datasets and LLMs demonstrate that models trained with PAFT exhibit strong robustness and generalization across a wide range of prompts, including unseen ones. This enhanced robustness improves both model performance and inference speed while maintaining training efficiency. Ablation studies further confirm the effectiveness of PAFT. 00footnotetext: Equal contribution, corresponding author.
@@
Chenxing Wei∗†§, Yao Shu∗§, Mingwen Ou∘, Ying Tiffany He†, Fei Richard Yu#†‡ †College of Computer Science and Software Engineering, Shenzhen University, China ∘Tsinghua Shenzhen International Graduate School, Tsinghua University, China §Guangdong Lab of AI and Digital Economy (SZ), China ‡School of Information Technology, Carleton University, Canada weichenxing2023@email.szu.edu.cn, shuyao@gml.ac.cn omm23@mails.tsinghua.edu.cn, richard.yu@ieee.org
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success across a diverse range of natural language processing (NLP) tasks Zhao et al. (2024); Xu et al. (2023). To further enhance the performance of LLMs on specific downstream tasks, supervised fine-tuning (SFT) has emerged as a widely adopted strategy Ouyang et al. (2022); Devlin et al. (2019). This approach typically involves augmenting input data with task-specific instructions and constructing dialogue datasets with expected outputs, enabling the model to effectively learn task-specific patterns during fine-tuning. Empirical studies have shown that SFT can substantially improve model performance on downstream tasks Raffel et al. (2023); Hu et al. (2023b); Wei et al. (2022). However, a critical limitation of this paradigm is its reliance on fixed instruction templates Mishra et al. (2022); Chung et al. (2022) for each downstream task. This rigidity often leads to overfitting, whereby models become excessively dependent on specific instruction patterns Zhang et al. (2024); Kung and Peng (2023). Consequently, during inference on downstream tasks, even minor deviations between user-provided instructions and the training instructions can result in significant performance degradation Mialon et al. (2023); Raman et al. (2023). This issue is particularly pronounced when LLM practitioners, who may lack domain expertise, provide prompts that deviate substantially from those used during SFT. In such scenarios, carefully fine-tuned models may experience drastic performance drops, occasionally approaching random guessing levels Voronov et al. (2024). Previous research has primarily focused on prompt tuning—introducing trainable vectors (soft prompts) to optimize performance Liu et al. (2022); Li and Liang (2021); Lester et al. (2021)—however, these methods inadvertently increase sensitivity to prompt variations Wen et al. (2023); Qin and Eisner (2021), resulting in significant performance fluctuations and increased costs associated with prompt engineering Han et al. (2024); Longpre et al. (2023). Prompt robustness in SFT has received limited attention, with most existing work focusing on in-context learning Zhu et al. (2024); Shi et al. (2024); Ishibashi et al. (2023).
@@
To address this critical gap, we present PAFT, an innovative fine-tuning framework designed to dynamically adapt to diverse prompts during training. To our knowledge, this is the first systematic approach to enhancing prompt robustness in SFT, a vital yet under-explored area. Unlike traditional methods, which often overfit to specific prompt patterns, PAFT enables models to grasp underlying task semantics, ensuring robust performance across various human-written prompts.As shown in Figure 2, PAFT operates in two phases: (1) Candidate Prompt Construction (Section 4.1) and (2) Dynamic Fine-Tuning (Section 4.2). Initially, a diverse set of high-quality synthetic prompts is generated, capturing essential task semantics while maintaining linguistic variability. During fine-tuning, a dynamic prompt sampling strategy is employed, randomly selecting prompts from our curated set to expose the model to a wide range of formulations. Extensive evaluations reveal that PAFT achieves three primary objectives: (1) significantly boosting model robustness and generalization across diverse prompts; (2) maintaining state-of-the-art performance on downstream tasks; and (3) potentially enhancing inference speed while preserving training efficiency. These findings indicate that PAFT represents a promising direction for developing more robust and user-friendly language models. Our key contributions are: (a) Through comprehensive experiments, we demonstrate that fine-tuning with fixed prompts significantly undermines the model’s robustness to prompt variations, leading to poor generalization on unseen prompts and severe performance degradation; and (b) We propose PAFT, comprising candidate prompt construction and dynamic fine-tuning, a novel approach to enhance the prompt robustness of fine-tuned models. This approach ensures consistent and robust performance across a variety of test prompts, including those not encountered during training.
@@ 2 Related Work
Prompt Optimization
Effective prompt engineering is crucial for maximizing LLM performance, motivating various optimization techniques Chang et al. (2024); Li (2023); Diao et al. (2023); Sun et al. (2022). Methods like INSTINCT Lin et al. (2024) utilize neural bandits and LLM embeddings for efficient prompt search, while ZOPO Hu et al. (2024) improves efficiency through localized search. BATprompt Shi et al. (2024) incorporates robustness considerations in in-context learning by leveraging natural language perturbations. However, these methods often suffer from prompt fragility, exhibiting high sensitivity to even minor prompt alterations, particularly after fine-tuning. This limits LLM generalization in real-world applications. Our work addresses this limitation by prioritizing robustness across diverse prompt formulations, rather than optimizing for a single prompt.
Supervised Fine-Tuning (SFT)
SFT is a dominant paradigm for adapting LLMs, valued for its efficiency. Two main SFT approaches exist: soft prompt tuning (optimizing continuous vectors prepended to the input while freezing base model parameters) Li and Liang (2021); Liu et al. (2022), and full/parameter-efficient fine-tuning (PEFT) Shu et al. (2024); Ouyang et al. (2022); Liu et al. (2021); Lester et al. (2021). Among PEFT techniques, Low-Rank Adaptation (LoRA) Hu et al. (2022) is widely used, freezing pre-trained parameters and introducing low-rank trainable matrices. Advanced LoRA variants further aim to mitigate overfitting and enhance generalization Chen et al. (2023); Si et al. (2024); Wei et al. (2024). However, these methods, while mitigating parameter-level overfitting, typically rely on fixed training prompts, thus neglecting prompt robustness. This is particularly problematic for soft prompt tuning, where models exhibit high sensitivity to prompt variations. Consequently, minor deviations from training prompts can drastically degrade performance. To address this, we propose PAFT, a novel framework that prioritizes prompt robustness while preserving computational advantages. By decoupling model performance from specific prompt formulations, PAFT significantly enhances the adaptability and reliability of fine-tuned models.
@@ 3 Preliminaries
To systematically study the impact of prompt variations on fine-tuned models, we use LoRA Hu et al. (2022) as an illustrative example and conduct comprehensive preliminary experiments on multiple downstream tasks to assess prompt sensitivity and robustness. These tasks include natural language inference, question answering, and reading comprehension, using the LLaMA3-8B Meta (2024) model. We constructed a comprehensive set of over 450 prompts, covering a wide range of language styles, task-specific instructions, and formatting variations. Figure 3 presents a statistical analysis of the accuracy distribution for both the base model and SFT model across these prompts, revealing a key finding: prompt selection significantly influences model performance, with considerable accuracy variation observed across prompts, irrespective of the downstream task. Only a small fraction (typically less than 10%) of prompts yields near-optimal performance; some even degrade accuracy to near-random levels. Minor prompt modifications (e.g., rephrasing, punctuation, reordering) induce substantial fluctuations. For example, the addition of "Question" improves accuracy by 20% (Figure 1). This sensitivity highlights the fragility of current fine-tuning methods and their strong dependence on specific prompt formulations. These findings align with prior work He et al. (2024); Voronov et al. (2024); Salinas and Morstatter (2024); Min et al. (2022); Gao et al. (2021); however, we demonstrate that this sensitivity persists across tasks, suggesting a fundamental limitation of current PEFT paradigms. Motivated by these findings, we propose PAFT, addressing prompt robustness by decoupling performance from specific formulations, ensuring consistent results across diverse prompts, and significantly enhancing the practical applicability of fine-tuned models in real-world scenarios where prompt variations are inevitable.
@@ 4 The PAFT Framework
To improve the prompt robustness of LLMs, we propose the PAFT framework in Figure 2. As shown in Figure 2, the PAFT framework consists of two key stages: candidate prompt construction (see Section 4.1 for details) and dynamic fine-tuning (see Section 4.2 for details).
@@ 4.1 Candidate Prompt Construction
To ensure the robustness and effectiveness of PAFT across diverse prompts, we design a comprehensive prompt construction framework that aims to generate diverse and meaningful candidate prompts efficiently, enabling the model to generalize across different prompt formats. Our approach leverages the powerful generative capabilities of LLMs Kohl et al. (2024) and comprises three key phases: First, recognizing the inherent variability in how different LLMs interpret downstream tasks due to variations in pre-training data, model architectures, and optimization objectives Minaee et al. (2024); Zhao et al. (2024), we employ a multi-model approach, selecting 10 mainstream LLMs according to their generation capabilities, including models from OpenAI et al. (2024); Bai et al. (2023); Ouyang et al. (2022), and other widely used commercial LLMs, for prompt generation. This diverse selection ensures broad coverage of potential prompt formulations, capturing variations in linguistic style, task interpretation, and instructional clarity, thereby mitigating biases towards any single model’s prompt generation tendencies. Second, we employ a dual-strategy approach, combining few-shot and zero-shot techniques to balance prompt quality and diversity. For few-shot prompting, we leverage principles from in-context learning, providing each LLM with carefully curated, human-crafted examples to guide the generation of semantically coherent and task-relevant prompts, ensuring meaningfulness and alignment with the intended task. For zero-shot prompting, we prioritize diversity by allowing LLMs to generate prompts without explicit examples, thus encouraging a wider range of linguistic styles, structural variations, and task formulations. Specifically, we generate 20 prompts using each strategy, resulting in a comprehensive set encompassing both high-quality prompts (derived from few-shot prompting) and diverse, potentially less optimal prompts (derived from zero-shot prompting). This balanced approach exposes the model to a realistic distribution of prompt quality during training, thereby enhancing its robustness to real-world scenarios where prompt quality may vary significantly. Finally, to rigorously evaluate the robustness of PAFT, we randomly partition the generated prompts into training and test sets using an 8:1 ratio. Crucially, the training and test sets contain entirely distinct prompts, ensuring evaluation on completely unseen formulations. This partitioning strategy enables the construction of training data that exposes the model to a wide range of prompt styles while providing a robust testbed for assessing generalization to novel prompts. By decoupling training and test prompts, we confirm that performance improvements reflect a genuine ability to handle diverse and unseen prompt formulations, rather than overfitting to specific prompt patterns. This comprehensive framework ensures that PAFT learns task semantics independently of specific prompt patterns, enabling effective generalization across a wide range of real-world scenarios, and provides a scalable and cost-effective solution for improving prompt robustness in LLMs.
@@ 4.2 Dynamic Fine-Tuning
The dynamic fine-tuning process in our PAFT framework is designed to enhance the robustness of LLMs to diverse prompt formulations while preserving high performance on downstream tasks. As illustrated in Algorithm 1, during each training epoch t, a prompt p is randomly sampled from a diverse set of synthetically generated candidate prompts (line 4 in Algorithm 1), ensuring exposure to a wide range of linguistic styles and task formulations. For each data point (line 6 in Algorithm 1), the selected prompt is reused for consecutive training steps (lines 7-9 in Algorithm 1), and the input is constructed by combining the prompt with the data point (line 7 in Algorithm 1). The model parameters are then updated using stochastic gradient-based optimization methods, such as SGD (Sra et al., 2011) or AdamW (Loshchilov and Hutter, 2019) (line 8 in Algorithm 1), enabling the model to learn task-specific semantics while adapting to the formulation of prompt. After every steps, a new prompt is sampled from to replace the current one (lines 10-11 in Algorithm 1), ensuring that the model is exposed to multiple prompts within a single epoch. At the end of each epoch, the model parameters are initialized with the final parameters from the previous epoch, (line 12 in Algorithm 1), ensuring continuity in the learning process. After epochs, the fine-tuned model parameters achieve consistent performance across a wide range of prompts (line 16 in Algorithm 1), including those not encountered during training. This makes PAFT particularly suitable for real-world applications where prompt quality and style may vary significantly, such as when users lack domain expertise or when prompts are generated automatically. By decoupling model performance from fixed prompt formulations, PAFT addresses a key limitation of traditional fine-tuning methods, ensuring robust performance without requiring extensive prompt engineering. The dynamic fine-tuning strategy enhances both the robustness and generalization of fine-tuned models while maintaining computational efficiency, making it a practical solution for improving the adaptability of LLMs in diverse settings.
@@
| user-specified prompt | 92.35 | 2.78 | 0% | 77.87 | 2.36 | 0% | 78.16 | 7.97 | 0% | 79.88 | 6.32 | 22% | 81.05 | 4.45 | 4% | 81.86 | 4.78 | 5% |
| TopAccuracy prompt | 91.27 | 2.79 | 86% | 75.96 | 3.89 | 0% | 66.77 | 3.94 | 0% | 84.81 | 4.06 | 59% | 82.45 | 3.26 | 14% | 80.25 | 3.63 | 32% |
@@
| PAFT Improvement | +1.37 | -1.09 | 14% | +5.81 | -1.11 | 73% | +3.93 | -3.00 | 100% | +2.45 | -1.83 | 29% | +2.72 | -1.55 | 38% | +4.25 | -1.94 | 54% |
@@
| user-specified prompt | 6.52 | 0.98 | 3.27 | 8.23 | 4.75 |
| TopAccuracy prompt | 5.75 | 1.13 | 2.76 | 7.56 | 4.30 |

### v2

@@ Abstract
Fine-tuning large language models (LLMs) often causes overfitting to specific prompt wording, where minor phrasing variations drastically reduce performance. To address this, we propose Prompt-Agnostic Fine-Tuning (PAFT), a method that enhances robustness through dynamic prompt variation during training. PAFT first generates diverse synthetic prompts, then continuously samples from this set to construct training instances, forcing models to learn fundamental task principles rather than surface-level patterns. Across systematic evaluations using both supervised fine-tuning (SFT) and reinforcement learning fine-tuning (RLFT), PAFT demonstrates substantially improved prompt robustness, achieving 7% higher generalization accuracy on unseen prompts than standard methods. In addition to enhanced robustness, PAFT consistently yields superior overall performance on established benchmarks for question answering, mathematical reasoning, and tool use. Notably, models trained with PAFT attain 3.2× faster inference speeds due to reduced prompt sensitivity. Ablation studies further validate effectiveness of PAFT, while theoretical analysis reveals that PAFT can effectively enhance the cross-domain generalization ability of LLM. 00footnotetext: corresponding author.
@@
Chenxing Wei†§, Mingwen Ou∘, Ying He#†, Yao Shu#≀, Fei Yu‡ †College of Computer Science and Software Engineering, Shenzhen University, China ∘Tsinghua Shenzhen International Graduate School, Tsinghua University, China §Guangdong Lab of AI and Digital Economy (SZ), China ≀Hong Kong University of Science and Technology (Guangzhou), China ‡School of Information Technology, Carleton University, Canada weichenxing2023@email.szu.edu.cn, yaoshu@hkust-gz.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success across diverse natural language processing (NLP) tasks Zhao et al. (2024); Xu et al. (2023). To further enhance the performance of LLMs on specific downstream tasks, supervised fine-tuning (SFT) Ouyang et al. (2022); Devlin et al. (2019) and reinforcement learning fine-tuning (RLFT) Wang et al. (2024); Wei et al. (2025) has emerged as a widely adopted strategy. These methods typically augment input data with task-specific instructions and construct dialogue datasets with expected outputs, enabling models to learn task-specific patterns. Empirical studies have shown that SFT and RLFT can substantially improve model performance on downstream tasks Raffel et al. (2023); Hu et al. (2023); Wei et al. (2022).
However, as shown in Figure 1, a critical limitation of current fine-tuning methods is their lack of prompt robustness, as further detailed in Sec. 3. Reliance on fixed instruction prompts Mishra et al. (2022); Chung et al. (2022) often leads to overfitting on specific prompts patterns Zhang et al. (2024); Kung and Peng (2023). Consequently, models become brittle: minor deviations between user and training prompts can significantly degrade inference performance Mialon et al. (2023); Raman et al. (2023). This brittleness manifests, for example, as substantial accuracy drops in QA tasks with altered prompt phrasing Wei et al. (2024), or as poor instruction following in chatbots and AI agents when commands deviate from those encountered during training Hong et al. (2024); Sahoo et al. (2025). Such sensitivity also raises fairness and reliability concerns in algorithmic comparisons Voronov et al. (2024). This vulnerability is particularly acute when users, unfamiliar with specific SFT prompt structures, provide highly divergent inputs, potentially causing fine-tuned models to perform near random guessing levels Polo et al. (2024). Notably, prompt robustness in SFT has received limited attention, with most existing work focusing on in-context learning and prompt tuning Shi et al. (2024); Ishibashi et al. (2023).
@@
To address this critical gap, we introduce PAFT, a novel framework that dynamically adapts to diverse training prompts. To our knowledge, PAFT is the first systematic approach to improving prompt robustness in both SFT and RLFT, a vital but underexplored area. Unlike traditional methods prone to overfitting specific prompt patterns, PAFT enables models to grasp underlying task semantics, ensuring robust performance across varied human-written prompts. PAFT operates in two phases (Figure 2): first, constructing a diverse set of high-quality synthetic prompts that capture essential task semantics with linguistic variability (Sec. 4.1); second, employing dynamic fine-tuning by sampling from this curated set to expose the model to various formulations (Sec. 4.2). Extensive evaluations demonstrate that PAFT significantly boosts model robustness and generalization to diverse prompts, maintains state-of-the-art downstream performance, and can potentially improve inference speed while preserving training efficiency. These findings highlight PAFT as a promising direction for developing more robust, user-friendly language models.
Our key contributions are as follows: (a) We highlight that fine-tuning with fixed prompts results in poor generalization to unseen prompts and severe performance degradation (Sec. 3). (b) We propose PAFT, a novel framework incorporating candidate prompt construction and dynamic fine-tuning, to enhance the prompt robustness of fine-tuned models (Sec. 4). (c) We empirically demonstrate the consistent and robust performance of PAFT across diverse downstream tasks, fine-tuning algorithms, and varied test prompts, including those unseen during training (Sec. 5). (d) We provide theoretical evidence that PAFT effectively enhances the cross-domain generalization of LLMs (Sec. 6).
@@ 2 Related Work
Prompt Optimization.
Prompt engineering critically influences LLM performance, driving numerous prompt optimization approaches Chang et al. (2024); Li (2023); Diao et al. (2023); Sun et al. (2022). Notable methods include INSTINCT Lin et al. (2024), which leverages neural network bandits with LLM embeddings for search efficiency. ZOPO Hu et al. (2024), which employs localized search strategies. BATprompt Shi et al. (2024), which integrates robustness through natural language perturbations. While these approaches excel at identifying single high-performance prompts, models fine-tuned on such prompts remain vulnerable to prompt variations. Our work, in contrast, addresses this limitation by simultaneously enhancing prompt robustness and optimizing performance across the entire prompt space rather than focusing on isolated optimal prompts.
Fine-tuning (FT).
SFT and RLFT constitute the main paradigms for adapting LLMs to downstream tasks, prized for their efficiency. These approaches split into two categories: soft prompt tuning, which optimizes continuous input vectors while preserving base model parameters Kong et al. (2025); Wu et al. (2024); Hu et al. (2024); Lin et al. (2024); Li and Liang (2021); Liu et al. (2022), and full/parameter efficient fine-tuning (PEFT) Shu et al. (2024); Ouyang et al. (2022); Liu et al. (2021); Lester et al. (2021). Among PEFT techniques, Low-Rank Adaptation (LoRA) Hu et al. (2022) predominates by freezing pre-trained weights while introducing trainable low-rank matrices, with recent variants enhancing generalization and reducing overfitting Chen et al. (2023); Si et al. (2024); Wei et al. (2024). Instruction tuning Sanh et al. (2022) further improves ability of model to follow diverse task-specific instructions. However, existing methods—particularly soft prompt tuning—still exhibit limited prompt robustness, leaving models vulnerable to prompt variations. Our work addresses this critical limitation while maintaining computational efficiency.
@@ 3 Preliminaries
To systematically study the impact of prompt variations on fine-tuned models, we conducted comprehensive experiments across multiple downstream tasks using LLaMA3-8B Meta (2024) with LoRA fine-tuning. We constructed a comprehensive set of over 450 prompts (both human-written and LLM-generated), covering a wide range of language styles, task-specific instructions, and formatting variations. Figure 3 presents a statistical analysis of the accuracy distribution for both the base and SFT models across these prompts, revealing a key finding: the formulation of the prompt dramatically influences the performance of the model regardless of the type of task, with only 10% of the prompts producing near-optimal results. Minor prompt modifications (e.g., rephrasing, punctuation, reordering) induce substantial fluctuations.
For example, the addition of "Question" improves accuracy by 20% (Figure 1). This sensitivity highlights the fragility of current fine-tuning methods and their strong dependence on specific prompt formulations. These findings align with prior work He et al. (2024); Voronov et al. (2024); Salinas and Morstatter (2024); Min et al. (2022); Gao et al. (2021b). This widespread sensitivity demonstrates a fundamental limitation in current fine-tuning approaches, extending findings from previous research across diverse task domains. Based on these insights, we propose PAFT, which decouples model performance from specific prompt formulations, ensuring consistent results across prompt variations and enhancing practical applicability in real-world scenarios.
@@ 4 The PAFT Framework
In this section, we introduce PAFT in detail. As shown in Figure 2, the PAFT framework consists of two key stages: candidate prompt construction (Sec. 4.1) and dynamic fine-tuning (Sec. 4.2).
@@ 4.1 Candidate Prompt Construction
To ensure the robustness and effectiveness of PAFT across diverse prompts, we design a comprehensive prompt construction framework that aims to generate diverse and meaningful candidate prompts efficiently, enabling the model to generalize across different prompt formats. Our approach leverages the powerful generative capabilities of LLMs Kohl et al. (2024) and comprises three key phases.
Diverse LLM Ensemble. We employ 10 mainstream LLMs with varied generation capabilities OpenAI et al. (2024); Bai et al. (2023); Ouyang et al. (2022) to capture the inherent variability in task interpretation stemming from differences in pre-training data, architectures, and optimization objectives Minaee et al. (2024); Zhao et al. (2024). This diversity ensures comprehensive coverage of prompt formulations across linguistic styles and instructional approaches, effectively mitigating single-model generation biases.
Dual Prompting Strategy. We combine few-shot and zero-shot techniques to balance quality and diversity. Few-shot prompting leverages in-context learning with curated human examples to generate task-aligned, semantically coherent prompts. Zero-shot prompting encourages diverse linguistic styles and structural variations without explicit examples. By generating 20 prompts with each strategy, we create a comprehensive set spanning high-quality and varied formulations, exposing the model to realistic prompt quality distributions and enhancing robustness to real-world scenarios. See Appendix D.1 for details.
Rigorous Evaluation Design. We randomly partition generated prompts into training and test sets (8:1 ratio), ensuring completely distinct prompts in each set. This approach exposes the model to diverse prompt styles during training while providing a robust testbed for assessing generalization to novel formulations, see Appendix D for more details of prompt sets. By evaluating on entirely unseen prompts, we confirm that performance improvements reflect genuine ability to handle diverse prompt formulations rather than overfitting to specific patterns. This framework ensures PAFT learns task semantics independently of prompt phrasing, enabling effective generalization across real-world scenarios while providing a scalable, cost-effective solution for improving prompt robustness.
@@ 4.2 Dynamic Fine-Tuning
Dynamic Fine-Tuning Algorithm. Our PAFT framework enhances the robustness of LLMs through systematic prompt diversification. As shown in Algorithm 1, each training epoch randomly samples a prompt from synthetic candidates (line 4), exposing the model to varied linguistic styles. For each data point (line 6), the selected prompt is reused for consecutive steps (lines 7-9), constructing inputs via (line 7) and updating parameters using gradient-based optimization like SGD (Sra et al., 2011) or AdamW (Loshchilov and Hutter, 2019) (line 8). After steps, a new prompt is sampled (lines 10-11), ensuring multiple prompt exposures per epoch. Each epoch initializes with final parameters from the previous one: (line 12), maintaining learning continuity until final parameters are achieved after epochs (line 16).
Benefits of Dynamic Fine-Tuning. Dynamic fine-tuning in PAFT significantly enhances LLM robustness and generalization by exposing the model to diverse prompts during training. This approach mitigates overfitting to fixed prompts, fostering the learning of more generalizable representations less sensitive to specific formulations. Consequently, PAFT achieves consistent performance across varied and unseen prompts, crucial for real-world applications with diverse user input. By reducing reliance on manual prompt engineering, dynamic fine-tuning offers an efficient and scalable solution for improving LLM adaptability.
@@
| User | 92.35 | 2.78 | 0% | 77.87 | 2.36 | 0% | 78.16 | 7.97 | 0% | 79.88 | 6.32 | 22% | 81.05 | 4.45 | 4% | 81.86 | 4.78 | 5% |
| TopAccuracy | 91.27 | 2.79 | 86% | 75.96 | 3.89 | 0% | 66.77 | 3.94 | 0% | 84.81 | 4.06 | 59% | 82.45 | 3.26 | 14% | 80.25 | 3.63 | 32% |
@@
| Improv. | +1.37 | -1.09 | 14% | +5.81 | -1.11 | 73% | +3.93 | -3.00 | 100% | +2.45 | -1.83 | 29% | +2.72 | -1.55 | 38% | +4.25 | -1.94 | 54% |
| Method | HumanEval | Xstory_cloze | Geometry3k | T-Eval | GSM8K |
| Base | 41.31 ( 10.36) | 48.23 ( 8.36) | 32.17 ( 15.36) | 58.97 ( 14.03) | 74.36 ( 21.37) |
| SFT | 49.63 ( 4.31) | 54.77 ( 4.79) | 37.94 ( 6.17) | 70.37 ( 8.14) | 81.47 ( 13.24) |
| PAFT | 54.24 ( 1.36) | 60.27 ( 0.73) | 40.19 ( 1.27) | 73.17 ( 3.27) | 85.71 ( 5.93) |
| Improv. | +4.61 (-2.95) | +5.50 (-4.06) | +2.25 (-4.90) | +2.80 (-4.87) | +4.24 (-7.31) |
@@
| User | 6.52 | 0.98 | 3.27 | 8.23 | 4.75 |
| TopAccuracy | 5.75 | 1.13 | 2.76 | 7.56 | 4.30 |